FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE E^{2}Π–X^{2}Σ^{+} TRANSITIONS OF SrH, SrD AND BaH.

\underline{R.S. RAM}, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; Department of Chemistry, University of Arizona, Tucson, AZ 85721, USA; K. TERESZCHUK, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; P.F. BERNATH, Department of Chemistry and Biochemistry, Old Dominion University, Norfolk, VA 23529 USA; Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; and K.A. WALKER, Department of Physics, University of Toronto, Toronto, Ont., M5S 1A7, Canada.

Emission spectra of SrH, SrD and BaH have been studied at high resolution using a Fourier transform spectrometer. The SrH and SrD molecules have been produced in a high temperature furnace by the reaction of strontium metal vapor with H_{2}/D_{2} in the presence of a slow flow of Ar gas. The spectra observed in the 18000–19000 cm^{−1} region consist of the 0–0 and 1–1 bands of the E^{2}Π–X^{2}Σ^{+} transition of the two isotopologues. A rotational analysis of these bands has been obtained by combining the present measurements with previously available pure rotation and vibration-rotation measurements for the ground state, and improved spectroscopic constants have been obtained for the E^{2}Π state.

The high resolution spectrum of the E^{2}Π–X^{2}Σ^{+} transition of BaH has been observed using a Ba hollow cathode lamp operated with 150 V and 300 mA current with a slow flow of 700 mTorr of Ar and 40 mTorr of H_{2} gases. The spectrum in the 14000–15600 cm^{−1} region was recorded at a resolution of 0.015 cm^{−1} with the 1-m Fourier transform spectrometer associated with the McMath-Pierce Telescope of the National Solar Observatory at Kitt Peak. A rotational analysis of the 0–0, 1–1 and 2–2 bands has been performed and improved spectroscopic constants were determined for the E^{2}Π state by combining the data with existing vibration-rotation measurements of the ground state.